OpenStudy (anonymous):

A homogeneous system of linear equations consists of six equations in five variables (i.e. unknowns). The rank of the coefficient matrix is 2. How many solutions does this system have?

OpenStudy (loser66):

infinite

OpenStudy (anonymous):

But the textbook says a homogenous linear system with more unknowns than equations has infinetly solutions. This one has 5 unknowns and 6 equations. It's less.

OpenStudy (loser66):

depend on the rank of augmented matrix. the rank of it is 2, that means you have 2 linearly independent equations, (just 2) and you have 5 variables. now, compare 2 and 5 which is bigger?

OpenStudy (anonymous):

My choices are: A.The system has no solution. B.The system has only the trivial solution. C.The system has a two-parameter family of solutions. D.The system has a three-parameter family of solutions. E.The system has a four-parameter family of solutions.

OpenStudy (loser66):

for example: x +y =5 2x +2y = 10 3x +3y = 15 the rank of the system is 1 although you have 3 equations and 2 unknowns. so, the solution is infinite set

OpenStudy (loser66):

in your particular case, 2 equations, 5 unknowns --> 3 free variables or 3 parameters Pick the choice, which one?? That's just terminology.

OpenStudy (anonymous):

So it's D

OpenStudy (loser66):

I said 2 equations . it means 2 linearly independent equations Yes, I think so . It's D

OpenStudy (anonymous):

Thank u!!!

OpenStudy (loser66):

np